# Thompson sampling for zero-inflated count outcomes with an application to the Drink Less mobile health study

### v1

@@ Abstract
Mobile health (mHealth) technologies aim to improve distal outcomes, such as clinical conditions, by optimizing proximal outcomes through just-in-time adaptive interventions. Contextual bandits provide a suitable framework for customizing such interventions according to individual time-varying contexts, intending to maximize cumulative proximal outcomes. However, unique challenges such as modeling count outcomes within bandit frameworks have hindered the widespread application of contextual bandits to mHealth studies. The current work addresses this challenge by leveraging count data models into online decision-making approaches. Specifically, we combine four common offline count data models (Poisson, negative binomial, zero-inflated Poisson, and zero-inflated negative binomial regressions) with Thompson sampling, a popular contextual bandit algorithm. The proposed algorithms are motivated by and evaluated on a real dataset from the Drink Less trial, where they are shown to improve user engagement with the mHealth system. The proposed methods are further evaluated on simulated data, achieving improvement in maximizing cumulative proximal outcomes over existing algorithms. Theoretical results on regret bounds are also derived. A user-friendly R package countts that implements the proposed methods for assessing contextual bandit algorithms is made publicly available at https://cran.r-project.org/web/packages/countts.
Keywords Count outcome  $\cdot$ Contextual bandits  $\cdot$ Mobile health  $\cdot$ Online decision-making  $\cdot$ Thompson Sampling
@@ 1 Introduction
Mobile health (mHealth) technologies use mobile devices to manage health and wellness by continuously observing contextual data such as location, step count, or other real-time individual information. The just-in-time adaptive intervention (JITAI) is a unique intervention design for delivering mHealth support, which can customize the content and delivery of support based on a user’s changing state and environment (Nahum-Shani et al., 2018). This intervention design can be conceptualized as an online decision-making problem, where the mobile phone delivers an intervention based on contextual information and monitors a proximal outcome at each decision time. The aim is to improve an ultimate distal outcome, such as reducing alcohol consumption, by maximizing the cumulative proximal outcome, such as engagement with an alcohol reduction app (Bell et al., 2020).
Contextual bandits provide an ideal framework for customizing JITAIs, where an agent continually observes users’ real-time contexts and selects from a range of actions or interventions based on available knowledge to maximize their cumulative proximal outcome or reward over a given time horizon (Tewari and Murphy, 2017). To achieve this, the agent must balance between exploiting the actions associated with the current highest expected outcomes and exploring other unexplored actions that may further improve the expected outcomes. (Dumitrascu et al., 2018).
The Thompson sampling (TS) approach, originally proposed by Thompson (1933) in the context of clinical trials, represents a Bayesian strategy for addressing the above exploration-exploitation trade-off. The underlying principle of TS is to select an action according to its posterior probability of being optimal, that is, leading to the highest expected outcome. Several empirical studies revealed that TS outperforms different bandit alternatives in various simulated and real data settings (Scott, 2010; Chapelle and Li, 2011). Recent theoretical research has also proved regret bounds for TS methods under linear bandit, generalized linear bandit, and full reinforcement learning settings (Agrawal and Goyal, 2013; Abeille and Lazaric, 2017; Russo and Van Roy, 2014, 2016). Moreover, TS-based algorithms have been used in mHealth studies, including DIAMANTE (Aguilera et al., 2020), HeartSteps II (Spruijt-Metz et al., 2022), and StratosPHere 2 (Deliu et al., 2023). Improved TS algorithms have also been developed to properly personalize JITAIs in mHealth settings (Liao et al., 2020; Tomkins et al., 2021; Kumar et al., 2023).
@@
In offline statistical analysis, several approaches or techniques have been developed to address these issues (Payne et al., 2017; Feng, 2021). Among the alternatives to Poisson regression, the negative binomial regression is frequently utilized, especially given its variance being a quadratic function of the mean, thereby facilitating the modeling of overdispersed count data (Ver Hoef and Boveng, 2007). In scenarios with excess zeros, the zero-inflated Poisson regression provides a robust framework. The model posits a latent mixture of a zero component and a Poisson component, i.e., with probability $p$ the only possible observation is 00, and with probability $1-p$, a Poisson random variable is observed (Lambert, 1992). To address additional overdispersion in the Poisson component, the zero-inflated negative binomial model is employed (Garay et al., 2011). Although these advancements have significantly improved the analysis of count data in offline settings, the adaptation and application of these models to online learning frameworks are not well studied.
@@
This work addresses the current lack of effective algorithms for contextual bandits with count outcomes in mHealth applications. One potential approach is to extend the existing generalized linear bandit framework to account for count outcomes (Filippi et al., 2010; Li et al., 2017; Kveton et al., 2020). However, these studies mainly focus on logistic bandits and lack discussion on count data. Moreover, many count regressions, such as negative binomial regression with an unknown dispersion parameter, zero-inflated Poisson regression, and zero-inflated negative binomial regression, do not belong to the generalized linear model framework (Cameron and Trivedi, 2013). Our current work aims to develop specific algorithms that are tailored to count data, such as the Poisson bandit or negative binomial bandit, which could address issues such as overdispersion and zero-inflation. Developing effective algorithms for count outcomes in contextual bandit settings could potentially improve the personalization of JITAIs and ultimately lead to better health outcomes.
@@
Both Bayesian and frequentist regret bounds have been derived to assess the theoretical performances of the proposed algorithms. To evaluate their empirical performances, we have conducted simulation studies under a wide range of scenarios. In addition, we have applied our approaches to the Drink Less data, obtained from a micro-randomized trial (MRT) designed to optimize the content and sequencing of push notifications in a behavior change app that helps people reduce alcohol consumption (Bell et al., 2020). Contextual bandits can enhance the utility of an MRT in several ways. For example, by adapting randomization probabilities in favor of the most effective interventions, they can offer a benefit to current users and directly translate clinical trial efficacy to real-world effectiveness by continuously learning from accumulating data.
@@
The structure of this article is as follows: Section 2 details our motivational study, the Drink Less MRT, while Section 3 outlines the problem setup and explores various count models. Section 4 introduces the TS-Count strategy, which integrates four count models, and discusses the approximation method for posterior sampling as well as the metric used to evaluate their performances. Theoretical results supporting our proposed algorithms are provided in Section 5. Section 6 assesses the performance of these algorithms via simulations, and Section 7 illustrates their application using the Drink Less MRT dataset. Finally, Section 8 concludes the article with a comprehensive discussion.
@@ 2 Motivating Example: the Drink Less Study
Our work is motivated by an mHealth app called Drink Less. This app targets UK adults drinking alcohol at increasing and higher risk levels in reducing hazardous and harmful alcohol consumption (Garnett et al., 2019, 2021).
The study is based on an MRT design: a useful experimental design that provides high-quality data to optimize the timing, content, and sequencing for mHealth interventions, intending to sustain engagement (Klasnja et al., 2015; Qian et al., 2022; Liu et al., 2023a). A 30-day MRT involving 349 participants was conducted to refine the push notification strategy for the Drink Less app (Bell et al., 2020, 2023). Participants with a baseline Alcohol Use Disorders Identification Test (AUDIT) score of 8 or higher (Bohn et al., 1995; Saunders et al., 1993), who resided in the UK, were at least 18 years old, and desired to drink less, were recruited into the trial. During the MRT, participants were randomly assigned each day at 8 p.m. to either receive a push notification (with a probability of 60%) or not (with a probability of 40%). However, low engagement levels can significantly decrease the app’s effectiveness. An exploratory proximal outcome was the overall number of screen views from 8 p.m. to 9 p.m., serving as a count-valued measure of the depth of engagement with the app. Figure 1 illustrates the empirical distribution and the corresponding Q-Q plot for this outcome, defined as the number of screen views recorded between 8 p.m. and 9 p.m. following the receipt of a push notification or no intervention. The analysis reveals that the proximal outcome deviates from a normal distribution, characterized notably by a high frequency of zero values. Additionally, the study gathered contextual data, including age, AUDIT score, gender, and occupation type, to explore potential moderating factors influencing the intervention’s effectiveness.
Motivated by this, the current article introduces a novel contextual bandit algorithm to optimize mHealth intervention delivery, focusing on count-based proximal outcomes. To assess the efficacy of the proposed approaches, we utilize data from the Drink Less MRT to build a simulated testing environment, as detailed in Section 7.
@@ 3.1 Notations and Problem Setup
We consider the setting where the learning algorithm embedded in a mobile app must decide whether to deliver an intervention and what type of intervention to provide to a given user over the follow-up time. The latter is defined on a time horizon $T$, with the decision occurring at discrete time points $t=1,2,...,T$. Let $\mathcal{A}$ denote the discrete set of available interventions at any time $t$ with cardinality $|\mathcal{A}|=K$. We denote the personal and contextual information of an individual at time $t$ by $X_{t}$. In the Drink Less study, for example, $X_{t}$ represents the real-time information of a user. After an intervention $A_{t}\in\mathcal{A}$ is delivered, we collect a stochastic proximal outcome $Y_{t}$. We assume that $Y_{t}$ is in count format, i.e., non-negative integers, with potential overdispersion and zero-inflation throughout the analysis. Further, let $\phi(A_{t},X_{t})\in\mathbb{R}^{d}$ denote the intervention-context feature, which includes the interactions between intervention and context.
As the environment for generating the proximal outcomes is unknown, the algorithm within the mobile app must estimate the expected outcomes for each intervention. This presents a trade-off between exploiting the best intervention so far versus exploring alternative interventions that may lead to increased expected outcomes. Contextual bandits offer a framework for addressing this trade-off in online decision-making problems (Tewari and Murphy, 2017). At each decision time $t$, contextual bandit algorithms utilize historical observations, say $D_{t-1}=\left\{X_{i},A_{i},Y_{i};i=1,\cdots,t-1\right\}$, to update the estimates for model parameters, e.g., the maximum likelihood estimation (MLE) method. A naive strategy can be the greedy algorithm (Sutton and Barto, 2018), which always exploits: at each decision time $t$, it selects the intervention with the highest expected proximal outcome. However, this can lead to unreliable algorithms that do not explore sufficiently and become stuck on suboptimal interventions. There are two alternative bandit approaches recognized for their ability to deal with this issue: the upper confidence bound method, which relies on “optimism in the face of uncertainty” (Filippi et al., 2010; Li et al., 2017), and the TS approach, which is based on the Bayesian framework (Russo et al., 2018). We adopt the TS strategy for two reasons: first, it has a valid theoretical guarantee and excellent empirical performance (Agrawal and Goyal, 2013; Abeille and Lazaric, 2017; Russo and Van Roy, 2014, 2016); second, it is a randomized strategy that assigns probabilities to each intervention option, making it suitable for MRTs and allowing for valid post-study causal inference (Liao et al., 2020).
@@ 3.2 Count Data Models: A Review
In this section, we provide a brief overview of two widely used regression models for count data, namely, Poisson regression and negative binomial regression. In addition, we also discuss two variations of these basic models, namely, zero-inflated Poisson regression and zero-inflated negative binomial regression, which are particularly useful in cases where there is potential zero-inflation.
@@ 3.2.1 Poisson Regression
It is also worth noting that the only condition for consistency is the correct specification of the conditional mean model (Gourieroux et al., 1984). However, due to the common issues with count data, i.e., overdispersion and zero-inflation, other models may be preferred over the basic Poisson regression.
@@ 3.2.2 Negative Binomial Regression
The negative binomial (NB) model is a popular alternative to handle overdispersion in count data analysis. The NB model is defined by the probability mass function
@@ 3.2.3 Zero-inflated Poisson Regression
Count data with extra zeros is common in many real-world applications. The zero-inflated Poisson (ZIP) model can handle such data by combining a distribution that is degenerate at zero with a Poisson distribution. In ZIP regression, the outcome $Y_{t}$ has a probability mass function given by
@@ 3.2.4 Zero-inflated Negative Binomial Regression
If the data suggests additional overdispersion after accounting for zero-inflation, it may be necessary to consider the zero-inflated negative binomial (ZINB) model that mixes a distribution degenerate at zero with a NB distribution. This is a more flexible approach compared to the ZIP model. In ZINB regression, $Y_{t}$ has a probability mass function given by
@@ 4 The TS-Count Strategy
In this section we present the TS-Count strategy for optimizing the delivery of mHealth interventions. The Thompson sampling strategy is closely connected to Bayesian inference, which select the optimal intervention according to its posterior probability of being optimal. Algorithm 1 presents the pseudocode for TS-Count, specialized to the case of a Poisson or NB bandit, while Algorithm 2 displays TS-Count in the context of a ZIP or ZINB bandit. Depending on the model used for $Y_{t}$, the strategy can be referred to as TS-Poisson, TS-NB, TS-ZIP, or TS-ZINB.
@@
2       Update the posterior distribution $Q_{t}(\beta)$ using data $D_{t-1}$ Observe context $X_{t}$ Sample $\tilde{\beta}_{t}$ from the updated posterior distribution $Q_{t}(\beta)$ Compute $A_{t}=\arg\max_{a\in\mathcal{A}}\exp\left\{\phi(a,X_{t})^{\top}\tilde{\beta}_{t}\right\}$ Offer intervention $A_{t}$ and observe the associated outcome $Y_{t}$

### v2

@@ Abstract
Mobile health (mHealth) interventions often aim to improve distal outcomes, such as clinical conditions, by optimizing proximal outcomes through just-in-time adaptive interventions. Contextual bandits provide a suitable framework for customizing such interventions according to individual time-varying contexts. However, unique challenges, such as modeling count outcomes within bandit frameworks, have hindered the widespread application of contextual bandits to mHealth studies. The current work addresses this challenge by leveraging count data models into online decision-making approaches. Specifically, we combine four common offline count data models (Poisson, negative binomial, zero-inflated Poisson, and zero-inflated negative binomial regressions) with Thompson sampling, a popular contextual bandit algorithm. The proposed algorithms are motivated by and evaluated on a real dataset from the Drink Less trial, where they are shown to improve user engagement with the mHealth platform. The proposed methods are further evaluated on simulated data, achieving improvement in maximizing cumulative proximal outcomes over existing algorithms. Theoretical results on regret bounds are also derived. The countts R package provides an implementation of our approach.
Count data,
Contextual bandits,
Mobile health,
Online decision-making,
Thompson sampling,
keywords:
[…]
, , , , and
@@ 1 Introduction
Mobile health (mHealth) interventions use mobile devices to manage health and wellness by continuously observing contextual data such as location, step count, or other real-time individual information. The just-in-time adaptive intervention (JITAI) is a unique intervention design for delivering mHealth support, which can customize the content and delivery of support based on a user’s changing state and environment (Nahum-Shani et al., 2018). This intervention design can be conceptualized as an online decision-making problem, where the mobile phone delivers an intervention based on contextual information and monitors a proximal outcome at each decision time. The aim is to improve an ultimate distal outcome, such as reducing alcohol consumption, by maximizing the cumulative proximal outcome, such as engagement with an alcohol reduction app (Bell et al., 2020).
Contextual bandits provide an ideal framework for customizing JITAIs, where an agent continually observes users’ real-time contexts and selects from a range of actions or interventions based on available knowledge to maximize their cumulative proximal outcome or reward over a given time horizon (Tewari and Murphy, 2017). To achieve this, the agent must balance between exploiting the actions associated with the current highest expected outcomes and exploring other unexplored actions that may further improve the expected outcomes (Dumitrascu, Feng and Engelhardt, 2018).
The Thompson sampling (TS) approach, originally proposed by Thompson (1933) in the context of clinical trials, represents a Bayesian strategy for addressing the above exploration-exploitation trade-off. The underlying principle of TS is to select an action according to its posterior probability of being optimal, that is, leading to the highest expected outcome. Several empirical studies revealed that TS outperforms different bandit alternatives in various simulated and real data settings (Scott, 2010; Chapelle and Li, 2011). Recent theoretical research has also proved regret bounds for TS methods under linear bandit, generalized linear bandit, and full reinforcement learning settings (Agrawal and Goyal, 2013; Abeille and Lazaric, 2017; Russo and Van Roy, 2014; Russo and Van Roy, 2016). Moreover, TS-based algorithms have been used in mHealth and clinical studies, including DIAMANTE (Aguilera et al., 2020), HeartSteps II (Spruijt-Metz et al., 2022), and StratosPHere 2 (Deliu et al., 2023). Improved TS algorithms have also been developed to properly personalize JITAIs in mHealth settings (Liao et al., 2020; Tomkins et al., 2021; Kumar et al., 2024; Greenewald et al., 2017; Huch et al., 2023).
@@
In offline statistical analysis, several approaches have been developed to address these issues (Payne et al., 2017; Feng, 2021). Among the alternatives to Poisson regression, the negative binomial (NB) regression is frequently utilized, especially given its variance being a quadratic function of the mean, thereby facilitating the modeling of overdispersed count data (Ver Hoef and Boveng, 2007). In scenarios with excess zeros, the zero-inflated Poisson (ZIP) regression provides a robust framework. This model posits a latent mixture of a zero component and a Poisson component, i.e., with probability $p$ the only possible observation is 00, and with probability $1-p$, a Poisson random variable is observed (Lambert, 1992). To address additional overdispersion in the Poisson component, the zero-inflated negative binomial (ZINB) model is employed (Garay et al., 2011). Although these advancements have significantly improved the analysis of count data in offline settings, the adaptation and application of these models to online learning frameworks are not well studied.
@@
The current paper aims to address the present lack of effective algorithms for contextual bandits with count outcomes in mHealth. One potential approach is to extend the existing generalized linear bandit framework to account for count outcomes (Filippi et al., 2010; Li, Lu and Zhou, 2017; Kveton et al., 2020). However, these studies mainly focus on logistic bandits and lack discussion on count data. Moreover, many count regressions, such as negative binomial regression with an unknown dispersion parameter, zero-inflated Poisson regression, and zero-inflated negative binomial regression, do not belong to the generalized linear model framework (Cameron and Trivedi, 2013). In the mHealth domain, the problem of zero inflation has been tackled in Trella et al. (2022, 2023), where a contextual Bayesian Poisson model is investigated in comparison to a Bayesian linear model (of a similar flavor to TS). However, their work focuses on the practical implementation of the algorithm with no insights on its theoretical properties and overdispersion is not taken into account.
@@
In this work, both Bayesian and frequentist regret bounds have been derived to assess the theoretical performances of the proposed algorithms. To evaluate their empirical performances, we have conducted simulation studies under a wide range of scenarios. In addition, we have applied our approaches to the Drink Less data, obtained from a micro-randomized trial (MRT) designed to optimize the content and sequencing of push notifications in a behavior change app that helps people reduce alcohol consumption (Bell et al., 2020). Contextual bandits can enhance the utility of an MRT in several ways. For example, by adapting randomization probabilities in favor of the most effective interventions, they can offer a benefit to current users and directly translate clinical trial efficacy to real-world effectiveness by continuously learning from accumulating data.
@@
The structure of this article is as follows. Section 2 details our motivating example, the Drink Less study, while Section 3 outlines the problem setup and explores various count models. Section 4 introduces the TS-Count strategy, which integrates four count models, and discusses the approximation method for posterior sampling as well as the metric used to evaluate their performances. Theoretical results supporting our proposed algorithms are provided in Section 5. Section 6 assesses the performance of these algorithms via simulations, and Section 7 illustrates their application using the Drink Less MRT dataset. Finally, Section 8 concludes the article with a comprehensive discussion.
@@ 2 Motivating Example: the Drink Less Study
Our work is motivated by a behavior change app called Drink Less. This app targets UK adults drinking alcohol at increasing and higher risk levels in reducing hazardous and harmful alcohol consumption (Garnett et al., 2019, 2021). The study is based on an MRT design: a useful experimental design that provides high-quality data to optimize the timing, content, and sequencing for mHealth interventions, intending to sustain engagement (Klasnja et al., 2015; Qian et al., 2022; Liu, Deliu and Chakraborty, 2023). A 30-day MRT involving 349 participants was conducted to refine the push notification strategy for the Drink Less app (Bell et al., 2020, 2023). Participants with a baseline Alcohol Use Disorders Identification Test (AUDIT) score of 8 or higher (Bohn, Babor and Kranzler, 1995; Saunders et al., 1993), who resided in the UK, were at least 18 years old, and desired to drink less, were recruited into the trial. During the MRT, participants were randomly assigned each day at 8 p.m. to either receive a push notification (with a probability of 60%) or not (with a probability of 40%). The study collected both baseline and time-varying contextual data, such as age, AUDIT score, gender, and occupation type. An exploratory proximal outcome was the overall number of screen views from 8 p.m. to 9 p.m., serving as a count measure of the depth of engagement with the app. Figure 1 illustrates the empirical distribution and the corresponding Q-Q plot for this outcome, defined as the number of screen views recorded between 8 p.m. and 9 p.m. following the receipt of a push notification or no intervention. The analysis reveals that the proximal outcome deviates from a normal distribution, characterized notably by a high frequency of zero values. Consequently, this article introduces innovative contextual bandit algorithms designed to optimize the delivery of interventions with count outcomes.
@@ 3.1 Notations and Problem Setup
We consider the setting where the learning algorithm embedded in a mobile app must decide whether to deliver an intervention and what type of intervention to provide to a given user over a predefined follow-up period. This period is defined on a time horizon $T$, with the decision occurring at discrete time points $t=1,2,...,T$. Let $\mathcal{A}$ denote the discrete set of available interventions at any time $t$ with cardinality $|\mathcal{A}|=K$. We denote the personal and contextual information of an individual at time $t$ by $X_{t}$. In the Drink Less study, for example, $X_{t}$ represents the real-time information of a user, typically selected based on domain knowledge. After an intervention $A_{t}\in\mathcal{A}$ is delivered, we collect a stochastic proximal outcome $Y_{t}$. We assume that $Y_{t}$ is count data, i.e., non-negative integers, with potential overdispersion and zero-inflation throughout the analysis. Further, let $\phi(A_{t},X_{t})\in\mathbb{R}^{d}$, with $d\geq 1$, denote the feature vector, which is a pre-specified function of the intervention $A_{t}$ and the context $X_{t}$. Typically, it includes the context, interventions, and interactions between intervention and context (see Section 7 for further illustration).
As the environment for generating the proximal outcomes is unknown, the algorithm within the mobile app must estimate the expected outcome for each intervention. This presents a trade-off between exploiting the best intervention so far versus exploring alternative interventions that might yield better results. Contextual bandits offer a framework for addressing this trade-off in online decision-making problems (Tewari and Murphy, 2017). At each decision time $t$, contextual bandit algorithms utilize historical observations, denoted as $D_{t}=\left\{X_{i},A_{i},Y_{i};i=1,\cdots,t-1\right\}$, to update the estimates of model parameters using approaches such as maximum likelihood estimation (MLE) method. A naive strategy is the greedy algorithm (Sutton and Barto, 2018), which always exploits by selecting the intervention with the highest estimated proximal outcome at each decision time $t$. However, this can lead to unreliable algorithms that do not explore sufficiently and get stuck at suboptimal interventions. There are two alternative bandit approaches recognized for their ability to deal with this issue: the upper confidence bound method, which relies on “optimism in the face of uncertainty” (Filippi et al., 2010; Li, Lu and Zhou, 2017), and the TS approach, rooted in the Bayesian framework (Russo et al., 2018). We have chosen to implement the TS strategy for two reasons: first, it has a valid theoretical guarantee and excellent empirical performance (Agrawal and Goyal, 2013; Abeille and Lazaric, 2017; Russo and Van Roy, 2014; Russo and Van Roy, 2016); second, it is a randomized strategy that assigns probabilities to each intervention option, making it suitable for MRTs and allowing for valid post-study causal inference (Liao et al., 2020; Liu et al., 2024).
@@ 3.2 Count Data Models: A Review
In this section, we provide a brief overview of two widely used regression models for count data, namely, Poisson regression and NB regression. In addition, we also discuss two variations of these basic models, namely, ZIP regression and ZINB regression, which are particularly useful in cases with zero-inflation.
@@ 3.2.1 Poisson Regression
It is also worth noting that the only condition for consistency is the correct specification of the conditional mean model (Gourieroux, Monfort and Trognon, 1984). However, due to the common issues with count data, i.e., overdispersion and zero-inflation, other models may be preferred over the basic Poisson regression.
@@ 3.2.2 Negative Binomial Regression
The NB model is a popular alternative to handle overdispersion in count data analysis. The NB model is defined by the probability mass function
@@ 3.2.3 Zero-inflated Poisson Regression
Count data with extra zeros is common in many real-world applications. The ZIP model can handle such data by combining a distribution that is degenerate at zero with a Poisson distribution. In ZIP regression, the outcome $Y_{t}$ has a probability mass function given by
@@ 3.2.4 Zero-inflated Negative Binomial Regression
Suppose the data suggests additional overdispersion after accounting for zero-inflation. In that case, it may be necessary to consider the ZINB model that mixes a distribution degenerate at zero with an NB distribution. This is a more flexible approach compared to the ZIP model. In ZINB regression, $Y_{t}$ has a probability mass function given by
@@ 4 The TS-Count Strategy
In this section, we present the TS-Count strategy for optimizing the delivery of mHealth interventions. The TS strategy is closely connected to Bayesian inference, which selects the optimal intervention according to its posterior probability of being optimal. Algorithm 1 presents the pseudocode for TS-Count, specialized to the case of a Poisson or NB bandit, while Algorithm 2 displays TS-Count in the context of a ZIP or ZINB bandit. Depending on the model used for $Y_{t}$, the strategy can be referred to as TS-Poisson, TS-NB, TS-ZIP, or TS-ZINB.
@@
2       Update the posterior distribution $Q_{t}(\beta)$ using data $D_{t}$
3      Observe context $X_{t}$
4      Sample $\tilde{\beta}_{t}$ from the updated posterior distribution $Q_{t}(\beta)$
5      Compute $A_{t}=\arg\max_{a\in\mathcal{A}}\exp\left\{\phi(a,X_{t})^{\top}\tilde{\beta}_{t}\right\}$
6      Offer intervention $A_{t}$ and observe the associated outcome $Y_{t}$
7 end for